# Provably Mitigating Overoptimization in RLHF: Your SFT Loss is Implicitly an Adversarial Regularizer

### v1

@@ Abstract
Aligning generative models with human preference via RLHF typically suffers from overoptimization, where an imperfectly learned reward model can misguide the generative model to output undesired responses. We investigate this problem in a principled manner by identifying the source of the misalignment as a form of distributional shift and uncertainty in learning human preferences. To mitigate overoptimization, we first propose a theoretical algorithm that chooses the best policy for an adversarially chosen reward model; one that simultaneously minimizes the maximum likelihood estimation of the loss and a reward penalty term. Here, the reward penalty term is introduced to prevent the policy from choosing actions with spurious high proxy rewards, resulting in provable sample efficiency of the algorithm under a partial coverage style condition. Moving from theory to practice, the proposed algorithm further enjoys an equivalent but surprisingly easy-to-implement reformulation. Using the equivalence between reward models and the corresponding optimal policy, the algorithm features a simple objective that combines: (i) a preference optimization loss that directly aligns the policy with human preference, and (ii) a supervised learning loss that explicitly imitates the policy with a (suitable) baseline distribution. In the context of aligning large language models (LLM), this objective fuses the direct preference optimization (DPO) loss with the supervised fune-tuning (SFT) loss to help mitigate the overoptimization towards undesired responses, for which we name the algorithm Regularized Preference Optimization (RPO). Experiments of aligning LLMs demonstrate the improved performance of RPO compared with DPO baselines. Our work sheds light on the interplay between preference optimization and SFT in tuning LLMs with both theoretical guarantees and empirical evidence.
@@ 1.1 Our Contributions
We highlight that the Imitation (SFT) loss serves as an important term to mitigate overoptimization. Even though the original DPO objective has already involved a KL regularization between the tuned LLM and the initial LLM, is not enough to prevent overoptimization. As elaborated in Section 4, the KL-regularization weight of the DPO objective could only control the scale of the gradient per training example, while the RPO objective can further modify the gradient direction. Calling back to the theoretical Algorithm 1, such a modification of gradient direction originates from the reward penalty in the adversarial objective for the reward model. This modification, as we expose in our theoretical analysis, helps to mitigate overoptimization. Thus, incoporating SFT loss in RLHF gives you a regularizer that provably mitigates overoptimization.
@@ 4 An Equivalent and Implementation-friendly Objective
$\widehat{r}=\mathop{\mathrm{argmin}}_{r\in\mathcal{R}}\bigg{\{}\eta\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}-\beta\cdot\log(\pi_{r}(a^{0}|x))\Big{]}+{\mathcal{L}}_{\mathcal{D}}\left(\beta\cdot\log\left(\frac{\pi_{r}(\cdot|\cdot)}{\pi^{\mathrm{ref}}(\cdot|\cdot)}\right)\right)\bigg{\}},$ (4.7)
@@
$\vspace{-2mm}\nabla_{\theta}\mathcal{L}_{\mathrm{RPO}}(\theta)$ $=\eta\beta\cdot\underbrace{\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}-\nabla_{\theta}\log(\pi_{\theta}(a^{0}|x))\Big{]}}_{\text{increase the alignment with the reference policy}}+\underbrace{\nabla_{\theta}{\mathcal{L}}_{\mathrm{DPO}}(\theta)}_{\text{decrease the DPO Loss}},$ (4.9)
@@ 5.1 Establishing the Sample Complexity of Maximin Objective (Algorithm 1)
$C_{\mu_{\mathcal{D}}}(\mathcal{R};\pi,\pi^{\mathrm{base}})\leq\left\|\frac{d_{0}\otimes\pi\otimes\pi^{\mathrm{base}}}{\mu_{\mathcal{D}}}\right\|_{\infty}=\sup_{(s,a^{1},a^{0})\in{\mathcal{S}}\times\mathcal{A}\times\mathcal{A}}\frac{d_{0}(x)\pi(a^{1}|x)\pi^{\mathrm{base}}(a^{0}|x)}{\mu_{\mathcal{D}}(s,a^{1},a^{0})}.$
@@
Please refer to Appendix A.1 for a detailed proof of Theorem 5.3. ∎
@@ 5.2 Equivalence between Maximin and Minimax Objectives
Under Assumption 5.5, it holds that (Lemma B.1)
@@
Please refer for Appendix B.1 for a detailed proof of Proposition 5.6. ∎
@@ 5.3 Generalization to New Prompt Distributions
Please refer to Appendix C for a detailed proof of Corollary 5.8. ∎
@@ 6 Experiments
According to Algorithm 2 and as we discussed in Remark 5.4, we implement RPO by adding an SFT loss (log probability of chosen responses in the preference dataset) to the original DPO loss. By comparing the evaluation performance on the test split of the training dataset, we select the hyperparameter $\eta$ as $0.005$ for both RPO (beta) and RPO (gemma). During the training of DPO and RPO, We keep the remaining hyperparameters including $\beta$, batch size, and learning rate to be the same for a fair comparison. Please see Appendix D.1 for a detailed training configuration.
@@
For the in-data distribution evaluation, we select the 200 prompts (which are not used in the selection of paremeter $\eta$) in the test split of the training dataset to let the reference model, DPO, and RPO generate the response respectively. We choose GPT-4 to annotate the preference in the response pairs. Though we instruct GPT-4 to give an annotation among win, lose, and tie (please see the full prompt in Appendix D.2), GPT-4 might still provide undesired annotations. Therefore, we filter all the undesired annotations and collect 150 examples for evaluation. We report the pairwise win rate among Ref., RPO, and DPO in Table 1 for both the beta and gemma series. To show a more illustrative comparison between DPO and RPO, we provide the barplot to report the number of pairwise examples annotated by GPT-4 in Figures 6 and 6. We observe that for both beta and gemma series, RPO has a better performance than DPO in terms of both RPO/DPO-SFT and RPO-DPO win rates. The performance improvement matches our theoretical results in Corollary 5.7, which shows the credit of the alleviation of overoptimization.
@@
We further evaluate the reference model, RPO-trained model, DPO-trained model, and the officially released DPO-trained model for both beta and gemma series in two standard LLM chat benchmarks: MT-bench (Zheng et al., 2024) and AlpacaEval 2.0 (Dubois et al., 2024). MT-Bench is a multi-turn benchmark that contains 160 questions across eight different domains of knowledge. The score for MT-Bench is evaluated by GPT-4 on a scale from 1 to 10. AlpacaEval 2.0 is a single-turn benchmark including 805 questions on different topics, mostly focused on helpfulness. The metrics of AlpacaEval 2.0 are the win rate and Length-Control (LC) win rate compared with GPT-4 Preview (11/06), where the annotator is also GPT-4 Preview (11/06) and LC win rate is proposed to mitigate the length bias of GPT-4. The results are summarized in Table 2, which shows that RPO consistently exceeds the performance of all the competitors (DPO, Reference model, and the officially released model trained by DPO) on MT-Bench and AlpacaEval 2.0. We also provide additional results on the pairwise win rate for these two benchmarks in Appendix D.3 to illustrate the performance improvement. The improvement can also be explained by the theoretical results in Corollary 5.8. Finally, we remark that RPO is a flexible plug-in module and can steadily improve the benchmark performance without changing the original training configuration or accessing extra preference data. This also sheds light on the potential of mitigating overoptimization for better alignment and generalization performance.
@@ Acknowledgement
The authors would like to thank Junyan Zhang for valuable discussions on the equivalence between the minimax and maxmin optimization.
Appendix A Proofs for Sample Complexity Analysis
A.1 Proof of Theorem 5.3
@@ Appendix A Proofs for Sample Complexity Analysis
$\mathrm{Gap}^{\pi}(\widehat{\pi})$ (A.1)
$\qquad=\mathbb{E}_{x\sim d_{0},a\sim\pi(\cdot|x)}\big{[}r^{\star}(x,a)\big{]}-\mathbb{E}_{x\sim d_{0},a\sim\widehat{\pi}(\cdot|x)}\big{[}r^{\star}(x,a)\big{]}$ (A.2)
$\qquad=\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{ref}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (A.3)
$\qquad\qquad-\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (A.4)
$\qquad\qquad+\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (A.5)
$\qquad\qquad-\mathbb{E}_{x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (A.6)
$\qquad\qquad+\beta\cdot\mathbb{E}_{x\sim d_{0}}\Big{[}\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}-\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (A.7)
@@
Term (A) $=\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (A.9)
$\qquad-\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\},$ (A.10)
@@
Term (B) $=\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (A.11)
$\qquad-\mathbb{E}_{x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),a^{0}\sim\pi^{\mathrm{base}(\cdot|x)}}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]},$ (A.12)

### v2

@@ Abstract
Aligning generative models with human preference via RLHF typically suffers from overoptimization, where an imperfectly learned reward model can misguide the generative model to output undesired responses. We investigate this problem in a principled manner by identifying the source of the misalignment as a form of distributional shift and uncertainty in learning human preferences. To mitigate overoptimization, we first propose a theoretical algorithm that chooses the best policy for an adversarially chosen reward model; one that simultaneously minimizes the maximum likelihood estimation of the loss and a reward penalty term. Here, the reward penalty term is introduced to prevent the policy from choosing actions with spurious high proxy rewards, resulting in provable sample efficiency of the algorithm under a partial coverage style condition. Moving from theory to practice, the proposed algorithm further enjoys an equivalent but surprisingly easy-to-implement reformulation. Using the equivalence between reward models and the corresponding optimal policy, the algorithm features a simple objective that combines: (i) a preference optimization loss that directly aligns the policy with human preference, and (ii) a supervised learning loss that explicitly imitates the policy with a (suitable) baseline distribution. In the context of aligning large language models (LLM), this objective fuses the direct preference optimization (DPO) loss with the supervised fine-tuning (SFT) loss to help mitigate the overoptimization towards undesired responses, for which we name the algorithm Regularized Preference Optimization (RPO). Experiments of aligning LLMs demonstrate the improved performance of RPO compared with DPO baselines. Our work sheds light on the interplay between preference optimization and SFT in tuning LLMs with both theoretical guarantees and empirical evidence.
@@ 1.1 Our Contributions
We highlight that the Imitation (SFT) loss serves as an important term to mitigate overoptimization. Even though the original DPO objective has already involved a KL regularization between the tuned LLM and the initial LLM, is not enough to prevent overoptimization. As elaborated in Section 4, the KL-regularization weight of the DPO objective could only control the scale of the gradient per training example, while the RPO objective can further modify the gradient direction. Calling back to the theoretical Algorithm 1, such a modification of gradient direction originates from the reward penalty in the adversarial objective for the reward model. This modification, as we expose in our theoretical analysis, helps to mitigate overoptimization. Thus, incorporating SFT loss in RLHF gives you a regularizer that provably mitigates overoptimization.
@@ 1.2 Related Works
Finally, we mention that the algorithm design of RPO is also related to the “pessimism” principle in the standard offline RL literature. It proposes to maintain a pessimistic estimate of the policy values or constrain the policy not to take unseen actions in the data to handle the challenge of the insufficient coverage of the dataset, e.g., (Jin et al., 2021; Uehara and Sun, 2021; Xie et al., 2021a, b; Rashidinejad et al., 2021; Yin and Wang, 2021; Xiong et al., 2022; Liu et al., 2022a; Shi et al., 2022; Zhan et al., 2022; Lu et al., 2022; Rashidinejad et al., 2022; Shi and Chi, 2024; Blanchet et al., 2024; Liu et al., 2022b; Liao et al., 2021; Fu et al., 2022). In contrast, we consider the offline RLHF problem and the techniques to obtain the objective of the RPO algorithm (see Section 4) along with its sample complexity analysis are new and different from these prior works on standard offline RL.
@@ 4 An Equivalent and Implementation-friendly Objective
$\widehat{r}=\mathop{\mathrm{argmin}}_{r\in\mathcal{R}}\bigg{\{}\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}-\beta\cdot\log(\pi_{r}(a^{0}|x))\Big{]}+{\mathcal{L}}_{\mathcal{D}}\left(\beta\cdot\log\left(\frac{\pi_{r}(\cdot|\cdot)}{\pi^{\mathrm{ref}}(\cdot|\cdot)}\right)\right)\bigg{\}},$ (4.7)
@@
$\vspace{-2mm}\nabla_{\theta}\mathcal{L}_{\mathrm{RPO}}(\theta)$ $=\eta\beta\cdot\underbrace{\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}-\nabla_{\theta}\log(\pi_{\theta}(a^{0}|x))\Big{]}}_{\text{increase the alignment with the baseline policy}}+\underbrace{\nabla_{\theta}{\mathcal{L}}_{\mathrm{DPO}}(\theta)}_{\text{decrease the DPO Loss}},$ (4.9)
@@ 5.1 Establishing the Sample Complexity of Maximin Objective (Algorithm 1)
$C_{\mu_{\mathcal{D}}}(\mathcal{R};\pi,\pi^{\mathrm{base}})\leq\left\|\frac{d_{0}\otimes\pi\otimes\pi^{\mathrm{base}}}{\mu_{\mathcal{D}}}\right\|_{\infty}=\sup_{(x,a^{1},a^{0})\in{\mathcal{S}}\times\mathcal{A}\times\mathcal{A}}\frac{d_{0}(x)\pi(a^{1}|x)\pi^{\mathrm{base}}(a^{0}|x)}{\mu_{\mathcal{D}}(x,a^{1},a^{0})}.$
@@
Please refer to Appendix B.1 for a detailed proof of Theorem 5.3. ∎
@@ 5.2 Equivalence between Maximin and Minimax Objectives
Under Assumption 5.5, it holds that (Lemma C.1)
@@
Please refer for Appendix C.1 for a detailed proof of Proposition 5.6. ∎
@@ 5.3 Generalization to New Prompt Distributions
Please refer to Appendix D for a detailed proof of Corollary 5.8. ∎
@@ 6 Experiments
According to Algorithm 2 and as we discussed in Remark 5.4, we implement RPO by adding an SFT loss (log probability of chosen responses in the preference dataset) to the original DPO loss. By comparing the evaluation performance on the test split of the training dataset, we select the hyperparameter $\eta$ as $0.005$ for both RPO (beta) and RPO (gemma). During the training of DPO and RPO, We keep the remaining hyperparameters including $\beta$, batch size, and learning rate to be the same for a fair comparison. Please see Appendix E.1 for a detailed training configuration.
@@
For the in-data distribution evaluation, we select the 200 prompts (which are not used in the selection of paremeter $\eta$) in the test split of the training dataset to let the reference model, DPO, and RPO generate the response respectively. We choose GPT-4 to annotate the preference in the response pairs. Though we instruct GPT-4 to give an annotation among win, lose, and tie (please see the full prompt in Appendix E.2), GPT-4 might still provide undesired annotations. Therefore, we filter all the undesired annotations and collect 150 examples for evaluation. We report the pairwise win rate among Ref., RPO, and DPO in Table 1 for both the beta and gemma series. To show a more illustrative comparison between DPO and RPO, we provide the barplot to report the number of pairwise examples annotated by GPT-4 in Figures 6 and 6. We observe that for both beta and gemma series, RPO has a better performance than DPO in terms of both RPO/DPO-SFT and RPO-DPO win rates. The performance improvement matches our theoretical results in Corollary 5.7, which shows the credit of the alleviation of overoptimization.
@@
We further evaluate the reference model, RPO-trained model, DPO-trained model, and the officially released DPO-trained model for both beta and gemma series in two standard LLM chat benchmarks: MT-bench (Zheng et al., 2024) and AlpacaEval 2.0 (Dubois et al., 2024). MT-Bench is a multi-turn benchmark that contains 160 questions across eight different domains of knowledge. The score for MT-Bench is evaluated by GPT-4 on a scale from 1 to 10. AlpacaEval 2.0 is a single-turn benchmark including 805 questions on different topics, mostly focused on helpfulness. The metrics of AlpacaEval 2.0 are the win rate and Length-Control (LC) win rate compared with GPT-4 Preview (11/06), where the annotator is also GPT-4 Preview (11/06) and LC win rate is proposed to mitigate the length bias of GPT-4. The results are summarized in Table 2, which shows that RPO consistently exceeds the performance of all the competitors (DPO, Reference model, and the officially released model trained by DPO) on MT-Bench and AlpacaEval 2.0. We also provide additional results on the pairwise win rate for these two benchmarks in Appendix E.3 to illustrate the performance improvement. The improvement can also be explained by the theoretical results in Corollary 5.8. Finally, we remark that RPO is a flexible plug-in module and can steadily improve the benchmark performance without changing the original training configuration or accessing extra preference data. This also sheds light on the potential of mitigating overoptimization for better alignment and generalization performance.
@@
RPO also improves math, reasoning, and coding abilities.
In addition to the MT-Bench and AlpacaEval 2.0 benchmarks, we further use benchmarks on math, reasoning, and coding tasks for evaluations of the RPO algorithm. Specifically, we choose the Grade School Math 8K (GSM8K) (Cobbe et al., 2021), AI2 Reasoning Challenge (ARC) (Clark et al., 2018), and Mostly Basic Python Programming (MBPP) (Austin et al., 2021) to measure the math, reasoning, and coding abilities of the model trained by RPO, respectively. We refer the readers to Appendix F for the detailed setups and results of these experiments.
@@ Acknowledgement
The authors would like to thank the anonymous reviewers for their helpful comments. The authors would also like to thank Junyan Zhang for valuable discussions on the equivalence between the minimax and maxmin optimization.
Appendix A Further Discussions
In this section, we make several further discussions on the RPO algorithm and its theory
Discussions on the partial coverage assumption (Assumption 5.2).
A sufficient condition to make this partial coverage condition (Assumption 5.2) hold is that the distribution of the offline dataset, which is $\mu_{\mathcal{D}}$, can well cover the joint distribution of $(a^{1},a^{0})\sim(\pi,\pi^{\mathrm{base}})$. Here to discuss focus on $\pi^{\mathrm{base}}=\pi^{\mathrm{chosen}}$ as we adopted in the experiment part.
First, we clarify that the offline dataset distribution $\mu_{\mathcal{D}}$ is not simply $(a^{1},a^{0})\sim(\pi^{\mathrm{unchosen}},\pi^{\mathrm{chosen}})$, since according to our definition (see Section 2) whether $a^{1}$ or $a^{0}$ is chosen is random and is determined by $y\in{0,1}$ obeying the BT model. Thus, $(a^{1},a^{0})\sim\mu_{\mathcal{D}}$ can be interpreted as a mixture of $(\pi^{\mathrm{unchosen}},\pi^{\mathrm{chosen}})$ and $(\pi^{\mathrm{chosen}},\pi^{\mathrm{unchosen}})$. This mixture probability would not be too small as long as the quality of $(a^{1},a^{0})$ does not vary too much, i.e., both of them are possible to be chosen, which is the case in practice. As a result, in the offline data distribution $(a^{1},a^{0})\sim\mu_{\mathcal{D}}$, both $a^{1}$ and $a^{0}$ partly comes from the chosen distribution $\pi^{\mathrm{chosen}}$.
Then in order for $\mu_{\mathcal{D}}$ to cover the joint distribution of $(a^{1},a^{0})\sim(\pi,\pi^{\mathrm{base}})$, it suffices to argue that $\pi^{\mathrm{chosen}}$ can cover the target policy $\pi$, which is then reduced back to the traditional coverage condition. Thus our assumption essentially requires that $\pi^{\mathrm{chosen}}$ well covers and only needs to cover the target policy $\pi$. This coincides with the spirit of the minimal data assumption in offline RL theory, i.e., the so-called partial coverage condition.
On the relationship between observed chosen probability and reward overoptimization.
First, we note that the actions and their chosen probabilities can be interpreted as a proxy of analyzing the underlying (estimated) reward model $\widehat{r}$ due to the representation $\pi_{\widehat{r}}(a|x)\propto\pi^{\mathrm{ref}}(a|x)\exp(\beta^{-1}\widehat{r}(x,a))$. Analyzing the (log) probabilities of the actions can be utilized to detect the mitigation of over-optimization, because according to the representation, an overestimated reward of a poor action would result in a higher probability of choosing this action, and would also cause a decay in the probability of choosing other better actions (since the probabilities are normalized to $1$).
To further showcase the ability of RPO to address overoptimization (through the lense of probability), consider the following theoretical example with only one state and three actions (Xu et al., 2024) where we can track everything clearly. It has three actions ${a,b,c}$ with $R^{\star}(a)=1,R^{\star}(b)=0.5,R^{\star}(c)=0$. The reference policy $\pi^{\mathrm{ref}}(a)=\pi^{\mathrm{ref}}(b)=0.4,\pi^{\mathrm{ref}}(c)=0.1$, and the dataset consists of one data point $\mathcal{D}=(a,b,1)$ (meaning action $a$ is preferred in the data). Then an ideally solved DPO objective would be $\pi_{\mathrm{DPO}}$ as long as $\pi^{\mathrm{DPO}}(b)=0$, and the value of $\pi^{\mathrm{DPO}}(a)$ can be arbitrarily chosen in $[0,1]$. Thus a possible solution to DPO would be $\pi^{\mathrm{DPO}}(a)=0.5,\pi^{\mathrm{DPO}}(b)=0$, and by the normalizing condition $\pi^{\mathrm{DPO}}(c)=0.5$, which is undesirable since the action $c$ has reward $R^{\star}(c)=0$. In contrast, solving the RPO objective would additionally require the maximization of $\pi_{\mathrm{RPO}}(a)$ due to the SFT regularization term, and thus the solution is shifted towards $\pi_{\mathrm{RPO}}(a)=1,\pi_{\mathrm{RPO}}(b)=\pi_{\mathrm{RPO}}(c)=0$, which is better than the DPO policy. Thus, RPO is able to prevent overoptimization towards poor actions that are less covered by the dataset (action $c$ here), therefore resulting in a better policy.
About the relationships and distinctions between PTX loss in (Taori et al., 2023) and the SFT loss of RPO.
The original PTX loss is an imitation loss calculated on the pretraining data. In contrast, the SFT loss in the RPO objective is an imitation loss calculated on the RLHF dataset. In more specific, our experiments use this SFT loss to imitate the chosen responses in the RLHF dataset. Thus the relationship is that they are both imitation loss which aims to mimic certain data distribution. The distinction is that they are calculated on different data sources. The SFT loss in the RPO objective naturally comes from our theoretical algorithm and provably serves as an important regularization term to mitigate overoptimization in offline RLHF.
About the computational complexity of the SFT loss gradient.
According to the paragraph Practical implementation in Section 6, RPO adds an additional SFT loss (the log probability of the chosen labels in the preference dataset) on the original DPO loss, where we highlight that the SFT loss is actually an intermediate quantity in the calculation of the DPO loss. Therefore, our proposed method does not incur any additional computation overhead compared with the vanilla DPO.
Appendix B Proofs for Sample Complexity Analysis
B.1 Proof of Theorem 5.3
@@ Appendix B Proofs for Sample Complexity Analysis
$\mathrm{Gap}^{\pi}(\widehat{\pi})$ (B.1)
$\qquad=\mathbb{E}_{x\sim d_{0},a\sim\pi(\cdot|x)}\big{[}r^{\star}(x,a)\big{]}-\mathbb{E}_{x\sim d_{0},a\sim\widehat{\pi}(\cdot|x)}\big{[}r^{\star}(x,a)\big{]}$ (B.2)
$\qquad=\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{ref}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.3)
$\qquad\qquad-\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (B.4)
$\qquad\qquad+\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (B.5)
$\qquad\qquad-\mathbb{E}_{x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.6)
$\qquad\qquad+\beta\cdot\mathbb{E}_{x\sim d_{0}}\Big{[}\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}-\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.7)
@@
Term (A) $=\mathbb{E}_{x\sim d_{0},a^{1}\sim\pi(\cdot|x),a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\pi(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}$ (B.9)
$\qquad-\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\},$ (B.10)
@@
Term (B) $=\eta^{-1}\cdot\min_{r\in\mathcal{R}}\left\{\eta\cdot\mathbb{E}_{\begin{subarray}{c}x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),\\ a^{0}\sim\pi^{\mathrm{base}}(\cdot|x)\end{subarray}}\Big{[}r(x,a^{1})-r(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]}+{\mathcal{L}}_{\mathcal{D}}(r)\right\}$ (B.11)
$\qquad-\mathbb{E}_{x\sim d_{0},a^{1}\sim\widehat{\pi}(\cdot|x),a^{0}\sim\pi^{\mathrm{base}(\cdot|x)}}\Big{[}r^{\star}(x,a^{1})-r^{\star}(x,a^{0})-\beta\cdot\mathrm{KL}\big{(}\widehat{\pi}(\cdot|x)\|\pi^{\mathrm{ref}}(\cdot|x)\big{)}\Big{]},$ (B.12)